# Breaking the Reviewer: Assessing the Vulnerability of Large Language Models in Automated Peer Review Under Textual Adversarial Attacks

### v1

@@ Abstract
Peer review is essential for maintaining academic quality, but the increasing volume of submissions places a significant burden on reviewers. Large language models (LLMs) offer potential assistance in this process, yet their susceptibility to textual adversarial attacks raises reliability concerns. This paper investigates the robustness of LLMs used as automated reviewers in the presence of such attacks. We focus on three key questions: (1) The effectiveness of LLMs in generating reviews compared to human reviewers. (2) The impact of adversarial attacks on the reliability of LLM-generated reviews. (3) Challenges and potential mitigation strategies for LLM-based review. Our evaluation reveals significant vulnerabilities, as text manipulations can distort LLM assessments. We offer a comprehensive evaluation of LLM performance in automated peer reviewing and analyze its robustness against adversarial attacks. Our findings emphasize the importance of addressing adversarial risks to ensure AI strengthens, rather than compromises, the integrity of scholarly communication.
@@
Tzu-Ling Lin1, Wei-Chih Chen1, Teng-Fang Hsiao1, Hou-I Liu1, Ya-Hsin Yeh1, Yu Kai Chan1, Wen-Sheng Lien1, Po-Yen Kuo1, Philip S. Yu2, Hong-Han Shuai1thanks: Corresponding author. 1National Yang Ming Chiao Tung University 2University of Illinois Chicago
@@ 1 Introduction
(1) How effectively can LLMs generate reviews that match the quality of human-written ones?
@@
To address these questions, we begin by evaluating the performance of leading LLMs, including closed-source models (e.g., GPT-4o, GPT-4o-mini), as well as open-source models (e.g., Llama-3.3-70B, Mistral-Small-3.1), on two core peer review tasks: review generation and score prediction. Our evaluation demonstrates the model’s capability to produce coherent and fluent reviews across multiple assessment aspects. We then apply common textual adversarial attack techniques to test the robustness of the LLM’s evaluations. Our findings reveal significant vulnerabilities, as the LLM’s assessments can be easily distorted through simple manipulations of the input text. Our contributions are threefold:
  • Aspect-based evaluation of LLM reviewer: We provide an in-depth analysis of an LLM’s performance in automatic paper reviewing tasks, comparing its outputs to those of human reviewers using aspect-based score.
  • Adversarial robustness analysis: We assess the impact of textual adversarial attacks on the reliability of LLM-generated reviews, highlighting specific vulnerabilities.
@@ 2.1 Automated Paper Review-Related Tasks
The surge in academic publications has overwhelmed traditional peer review systems Weissgerber et al. (2021); Severin and Chataway (2021); Shah (2022), prompting the development of automated tools for tasks like readability assessment, plagiarism detection, and statistical error checking Kilicoglu et al. (2018); Riedel et al. (2020, 2022); Zhang (2010); Nuijten et al. (2016); Checco et al. (2021); Schulz et al. (2022). While helpful, these tools lack the ability to evaluate scientific content or methodological soundness.
Recent advances in NLP enable deeper automation. Nikiforovskaya et al. (2020) and Yuan et al. (2022) proposed automatic review generation pipelines using BERT and BART, respectively. Lu et al. (2024) introduced a framework for LLM-driven scientific research and evaluation. Du et al. (2024) investigated LLMs as (meta-)reviewers in NLP, while Yu et al. (2024a) proposed a framework combining review generation and self-critique.
Several works Liu and Shah (2023); Robertson (2023); Zhou et al. (2024); Liang et al. (2024) assess GPT-4’s review quality, revealing its potential and limitations. Other studies Wu et al. (2025, 2024); Yu et al. (2024b) address the detectability of LLM-generated reviews, emphasizing the difficulty of avoiding false positives and negatives, which complicates efforts to safeguard review integrity.
2.2 Vulnerabilities of LLMs to Adversarial Text Attacks
Adversarial text attacks manipulate inputs at different levels—characters Gao et al. (2018); Ebrahimi et al. (2018); Belinkov and Bisk (2017); Wei et al. (2024), words Jin et al. (2020); Li et al. (2020); Maheshwary et al. (2021), or sentences Qi et al. (2021a, b)—to mislead models without altering the core meaning. These attacks, whether white-box or black-box, expose vulnerabilities in LLMs with minimal and often imperceptible changes.
Recent surveys Yao et al. (2024); Kumar (2024) highlight growing concerns, including data poisoning, prompt injection, and biases such as position and verbosity bias Liu et al. (2024); Saito et al. (2023), or self-enhancement bias Zheng et al. (2023). In peer review, Robertson (2023) and Raina et al. (2024) show that subtle adversarial edits can mislead GPT-based reviewers, inflating scores or obscuring flaws, underscoring the risks in relying solely on LLMs for evaluation tasks.
@@ 3 LLM Vulnerability on Paper Reviews
Figure 1: The scenario illustration for LLM under adversarial attack, which author can be potential attacks introduce specific patterns (e.g., typos) into paper content. Thus, the feedback generated by LLM reviewers was misleading.
In response to the growing use of LLMs in peer review Zhou et al. (2024); Tyser et al. (2024); Yu et al. (2024b); Lu et al. (2024), we investigate their ability to automatically generate scientific reviews and their robustness against textual adversarial attacks—subtle edits that preserve meaning but may influence model outputs.
@@ 3.1 Peer Review Task Formulation
Task 1: Review Generation. The most direct application of leveraging LLMs as reviewers is the automatic generation of reviews for academic papers. In this task, we investigate the capability of LLM to produce constructive peer reviews that are coherent and consistent with human evaluations. Building on the aspect labels provided by Yuan et al. (2022), we construct a predefined aspect typology aligned with the ACL review guidelines. This typology includes the categories: “Summary”, “Motivation”, “Substance”, “Originality”, “Soundness”, “Clarity”, “Replicability”, “Meaningful comparison”, and “None”. Except for “Summary”, each aspect is further classified into positive and negative sentiments to capture both strengths and weaknesses of the paper. The “None” tag is assigned to sentences that do not fit any categories.
@@
Task 2: Score Prediction. Score prediction serves as a vital complement to review generation by offering quantitative evaluations of a paper across multiple aspects Kang et al. (2018); Zhou et al. (2024). Its structured and numerical format provides clearer evaluation standards and enables more consistent comparisons. In this task, we assess the capability of LLMs to assign scores from different perspectives, simulating the nuanced judgments typically made by human reviewers.
@@
To provide additional insight into the model’s reasoning, we ask it to include a brief justification for each score (as illustrated in Appendix C). This practice mirrors that of human reviewers, who often accompany their ratings with explanatory comments. By comparing the LLM-generated scores and justifications with those provided by human reviewers, we can assess the alignment between automated and expert evaluations. Furthermore, this task enables us to study the impact of adversarial text perturbations on the model’s scoring behavior. By introducing subtle modifications to the input, we can observe how the LLM’s assigned scores shift across different aspects. This analysis reveals the model’s sensitivity to input variation and its robustness in maintaining consistent evaluations under potentially manipulated conditions.
@@ 3.2 Attack Focus Localization
In paper reviewing tasks, the content under review is often too lengthy, making it computationally infeasible to directly search for adversarial patterns across the entire paper. To address this challenge, we first identify the content areas of the paper that attract the most attention to the LLM while generating reviews by feeding the clean paper content into the model and requesting a review.
@@
=$=\{m_{1},m_{2},\dots,m_{N}\},$
where $\mathcal{M}$ is the modifiable sequences set for a single paper, $f_{r}(\cdot)$ denotes the review generation by LLM $f$ based on the input content. LCS identifies the longest common sequence matching algorithm.  [^1] LCS can be character-level or word-level depending on the attack method; implementation details could be found in Appendix D., $m$ represent modifiable subsequences, and $N$ is the total number of matched subsequences.
@@
, where $\mathcal{P}_{m}$ defines the possible set of perturbations for every modifiable sequence $m$, identified through the transformations specified by the attack method $A$. $\delta$ represents the textual perturbation applied to the original content, and $x^{*}_{adv}$ is the optimal perturbed content found through the search method of the attack. $f_{s}(\cdot)$ represents the sum of the score for all aspects predicted by LLM $f$.
@@
Total Score $\text{Total Score Shift}=\sum^{|\mathcal{D}|}\left(f_{s}(x^{*}_{adv})-f_{s}(x_{clean})\right)$
@@ 3.3 Textual Adversarial Attacks
Here, we introduce several textual adversarial attacks to explore the robustness of the review generation and score prediction procedure. We categorize adversarial attack methods based on common strategies in textual attacks. For each strategy, we introduce widely recognized representative approaches in black-box settings to test the limits of the robustness of the LLM.
@@
Word-Level Attacks. For word-level adversarial attacks, we implement TextFooler Jin et al. (2020) and BERT-Attack Li et al. (2020) to deceive the LLM. Both attacks focus on replacing words with their synonyms or similar terms in contextualized embedding space, maintaining semantic similarity while altering the input text.
Sentence-Level Attacks. For the sentence-level attacks, we use StyleAdv Qi et al. (2021a) as representatives for sentence-level attacks. These methods involve transforming sentences in the paper content into specific text styles and modifying stylistic elements without altering the semantic meaning. Such transformation are designed to be imperceptible to human readers but can potentially mislead the LLM.
To conclude, applying these adversarial attack methods enables us to assess the robustness of LLMs in automated peer review. By introducing perturbations at the character, word, and sentence levels, we identify vulnerabilities that compromise the LLM’s reliability. The examples of each attack are provided in Table 5, which offers a visualizations of how these perturbations alter the original sentences. Also, the algorithm of textual adversarial attacks can be found in Appendix E.
@@ 4.1 Experiment Settings
LLM Selection and Configuration. We adopt both closed- and open-source LLMs for automated peer review. From the closed-source family, we select GPT-4o and GPT-4o-mini due to their strong performance in text evaluation tasks Robertson (2023); Zhou et al. (2024), with the latter offering a cost-efficient alternative. For open-source models, we include LLaMA-3.3-70B and Mistral-small-3.1. Llama-3.3-70B is chosen for its strong general language capabilities, while Mistral-Small-3.1 has shown competitive performance with GPT-4o-mini, particularly in long-context understanding. Prompt templates and configurations are provided in Appendix A and B.
Dataset. PeerRead dataset Kang et al. (2018) is one of the earliest released datasets of peer reviews from the NLP community. Following Zhou et al. (2024), we select the ICLR-2017 subset for both review generation and score prediction tasks, which contains 427 papers (177 accepted and 255 rejected) with 1,304 reviews written by experts and aspect-specific scores assigned by the annotators. (The experiment result of Adversarial Attack Performance and Score Shift Comparison is provided in Appendix G)
AgentReview Jin et al. (2024) presents a large-scale dataset curated through their LLM-based peer review simulation framework. The dataset comprises papers from ICLR over four years (2020–2023), sourced from OpenReview, and includes 523 papers (350 rejected, 125 poster, 29 spotlight, and 19 oral) with 10,460 generated reviews on them.
Since the reviews in the both dataset are not manually annotated with aspect, we leverage the aspect tagger pre-trained on ASAP dataset Yuan et al. (2022) to automatically provide aspect tags for human-written reviews in the PeerRead dataset.
@@ 4.2 Evaluation Metrics
Review Quality Metrics. To address the question - How effective are LLMs in generating coherent and insightful reviews compared to human reviewers? We compare LLM-generated reviews with human-annotated reviews, treating the latter as ground truth. Following the metrics proposed by Yuan et al. (2022), we evaluate the text quality of the automated reviews across multiple dimensions. We aim to assess both the overall quality of the LLM-generated reviews and their consistency with human evaluations.
(1) Comprehensiveness: For automatically generated reviews, we expect them can cover diverse aspects of the paper. We calculate the Aspect Coverage (ACOV) metric to assess the generated review. In particular, the number of aspects in a predetermined aspect typology that a review has covered can be used to quantify aspect coverage.
$ACOV=\frac{\text{\# of aspect tags in generated review}}{\text{\# of predefined aspects}}$
(2) Similarity to human: To evaluate the semantic equivalence between LLM-generated reviews and multiple human-annotated ground truth reviews, we utilize classical statistical methods: ROUGE-1, ROUGE-2, ROUGE-L Lin (2004), and the BertScore metric Zhang et al. (2019). Rather than averaging scores across all human references, we select the maximum score between the LLM-generated review and the human-written ground truths. This approach assumes that closely aligning with every reference is unnecessary; instead, identifying the best match offers a more meaningful assessment of semantic equivalence.
Notably, bias and omissions may definitely exist in manually annotated reviews/scores. However, the human-written evaluation can still be the gold reference to a certain extent.

### v3

@@ Abstract
Peer review is essential for maintaining academic quality, but the increasing volume of submissions places a significant burden on reviewers. Large language models (LLMs) offer potential assistance in this process, yet their susceptibility to textual adversarial attacks raises reliability concerns. This paper investigates the robustness of LLMs used as automated reviewers in the presence of such attacks. We focus on three key questions: (1) The effectiveness of LLMs in generating reviews compared to human reviewers. (2) The impact of adversarial attacks on the reliability of LLM-generated reviews. (3) Challenges and potential mitigation strategies for LLM-based review. Our evaluation reveals significant vulnerabilities, as text manipulations can distort LLM assessments. We offer a comprehensive evaluation of LLM performance in automated peer reviewing and analyze its robustness against adversarial attacks. Our findings emphasize the importance of addressing adversarial risks to ensure AI strengthens, rather than compromises, the integrity of scholarly communication. Our project page can be found in https://github.com/Lin-TzuLing/Breaking-the-Reviewer.git
@@
Tzu-Ling Lin1, Wei-Chih Chen1, Teng-Fang Hsiao1, Hou-I Liu1, Ya-Hsin Yeh1, Yu-Kai Chan1, Wen-Sheng Lien1, Po-Yen Kuo1, Philip S. Yu2, Hong-Han Shuai1thanks: Corresponding author. 1National Yang Ming Chiao Tung University 2University of Illinois Chicago
@@ 1 Introduction
(1) How effectively can LLMs generate reviews that approach the quality of human-written ones?
@@
To address these questions, we begin by evaluating the performance of leading LLMs, including closed-source models (e.g., GPT-4o, GPT-4o-mini), as well as open-source models (e.g., Llama-3.3-70B, Mistral-small-3.1), on two core peer review tasks: review generation and score prediction. Our evaluation demonstrates the models’ capability to produce coherent and fluent reviews across multiple assessment aspects. We then apply common textual adversarial attack techniques to test the robustness of the LLM’s evaluations. Our findings reveal significant vulnerabilities, as the LLM’s assessments can be easily distorted through simple manipulations of the input text. Our contributions are threefold:
  • Aspect-based evaluation of LLM reviewer: We conduct an in-depth evaluation of LLM performance in aspect-based automatic paper reviewing, comparing outputs with human reviewers in the review generation task.
  • Adversarial robustness analysis: We investigate the reliability of LLM-generated reviews against textual adversarial attacks in the score prediction task and propose a novel method, Attack Focus Localization, to pinpoint vulnerabilities in lengthy academic documents.
@@ 2.1 Automated Paper Review-Related Tasks
The exponential growth in academic publications has significantly increased peer reviewers’ workload, straining traditional peer review systems Weissgerber et al. (2021); Severin and Chataway (2021); Shah (2022). To mitigate this burden, research in automated peer review has been developed and can be broadly divided into two streams: (i) pre-peer review screening tools and (ii) approaches leveraging advanced natural language processing (NLP) and large language models (LLMs).
Pre-peer review automation mainly focused on screening tasks that help ensure compliance with journal policies, such as readability assessment, plagiarism detection, and statistical error checking Kilicoglu et al. (2018); Riedel et al. (2020, 2022); Zhang (2010); Nuijten et al. (2016); Checco et al. (2021); Schulz et al. (2022). While effective in improving editorial efficiency, these tools are limited to surface-level checks and cannot evaluate a paper within its broader scientific context or assess methodological appropriateness.
The transition to NLP-based review generation reflects an effort to move beyond rule-based checks Kuznetsov et al. (2024). For example, Nikiforovskaya et al. (2020) integrated co-citation analysis with BERT-based summarization to generate review content, while Yuan et al. (2022) developed a system using BART and established criteria for assessing review quality. Compared to earlier screening tools, these approaches move closer to approximating human-like judgment, but they remain constrained by domain specificity, dataset scale, and the reliability of generated evaluations.
Building on these early NLP-based pipelines, recent research has turned to large language models (LLMs). Lu et al. (2024) introduced an LLM-driven scientific research and evaluation framework, Du et al. (2024) investigated LLMs as (meta-)reviewers in NLP venues, and Yu et al. (2024a) proposed a framework combining review generation with self-critique. The emergence of GPT-4 further prompted several studies Liu and Shah (2023); Robertson (2023); Zhou et al. (2024); Liang et al. (2024) that benchmarked LLM-generated reviews against human reviewers, highlighting both strengths (e.g., consistency, speed) and limitations (e.g., lack of in-depth domain expertise). In parallel, other works have examined the detectability of LLM-generated reviews Wu et al. (2025, 2024); Yu et al. (2024b), emphasizing the difficulty of avoiding false positives and negatives. Unlike earlier screening tools, these studies underscore the dual problem of harnessing LLMs for review assistance while safeguarding review integrity.
2.2 Vulnerabilities of Large Language Models to Adversarial Text Attacks
Adversarial text attacks subtly modify input content to mislead language models, resulting in altered outputs. Such attacks can be categorized along two dimensions: the granularity of perturbations and the attacker’s knowledge of the target model. At the perturbation level, character-level manipulations (insertion, deletion, swapping) can cause significant disruptions with minimal changes Gao et al. (2018); Ebrahimi et al. (2018); Belinkov and Bisk (2017), but generate nonsensical tokens and cause obvious typos. Word-level attacks, typically based on synonym replacement or contextual tweaks, preserve grammaticality yet can be easier to detect through lexical constraints Jin et al. (2020); Li et al. (2020); Maheshwary et al. (2021). Sentence-level attacks leverage paraphrasing or syntactic restructuring, which can be highly deceptive though computationally intensive Qi et al. (2021a, b). Attacks are also distinguished by attacker knowledge: white-box attacks exploit full model access, enabling gradient-based perturbations, whereas black-box attacks operate without internal information, trading precision for broader applicability in real-world scenarios.
While these techniques have long been studied in NLP, the rise of LLMs introduces new vulnerabilities. Yao et al. (2024); Kumar (2024) highlight threats such as data poisoning, backdoor attacks, and prompt injection. Compared to traditional models, LLMs also exhibit behavioral weaknesses: position and verbosity biases, where models favor earlier options or longer responses Liu et al. (2024); Saito et al. (2023), as well as self-enhancement bias, where models tend to portray themselves positively Zheng et al. (2023). These biases complicate evaluation, making LLMs vulnerable to adversarial inputs while simultaneously undermining their reliability as evaluators of their own outputs.
The impact of such vulnerabilities becomes particularly salient in high-stakes applications. In the context of peer review, Robertson (2023) observed that GPT-4 could detect inconsistencies after introducing adversarial modifications to abstracts, but it struggled with subtler manipulations such as shifts in formality. Similarly, Raina et al. (2024) showed that adversarial manipulation could significantly inflate evaluation scores, raising concerns about the fairness and reliability of automated LLM-based reviewing processes.
@@ 3 LLM Vulnerability on Paper Reviews
Figure 1: Illustration of a scenario where an LLM is under adversarial attack: the author may introduce specific patterns (e.g., typos) into the paper, causing the feedback generated by LLM reviewers to be misled.
In response to the growing use of LLMs in peer review Zhou et al. (2024); Tyser et al. (2024); Yu et al. (2024b); Lu et al. (2024), we investigate their ability to automatically generate scientific reviews and examine their robustness against textual adversarial attacks, which are subtle edits that preserve meaning but may influence model outputs.
@@ 3.1 Peer Review Task Formulation
Task 1: Review Generation. The most direct application of leveraging LLMs as reviewers is the automatic generation of reviews for academic papers. In this task, we investigate the capability of LLMs to produce constructive peer reviews that are coherent and consistent with human evaluations. Building on the aspect labels provided by Yuan et al. (2022), we construct a predefined aspect typology aligned with the ACL review guidelines. This typology includes the categories: “Summary”, “Motivation”, “Substance”, “Originality”, “Soundness”, “Clarity”, “Replicability”, “Meaningful comparison”, and “None”. Except for “Summary”, each aspect is further classified into positive and negative sentiments to capture both strengths and weaknesses of the paper. The “None” tag is assigned to sentences that do not fit any categories.
@@
Task 2: Score Prediction. Score prediction serves as a vital complement to review generation by offering quantitative evaluations of a paper across multiple aspects Kang et al. (2018); Zhou et al. (2024). Its structured and numerical format provides clearer evaluation standards and enables more consistent comparisons. In this task, we assess the capability of LLMs to assign scores from different perspectives, simulating the expert judgments typically made by human reviewers.
@@
To provide additional insight into the model’s reasoning, we ask it to include a brief justification for each score (as illustrated in Appendix A). This practice mirrors that of human reviewers, who often accompany their ratings with explanatory comments. By comparing the LLM-generated scores and justifications with those provided by human reviewers, we can assess the alignment between automated and expert evaluations. Furthermore, this task enables us to study the impact of adversarial text perturbations on the model’s scoring behavior. By introducing subtle modifications to the input, we can observe how the LLM’s assigned scores shift across different aspects. This analysis reveals the model’s sensitivity to input variation and its robustness in maintaining consistent evaluations under potentially manipulated conditions.
@@ 3.2 Attack Focus Localization
In paper reviewing tasks, the content under review is often too lengthy, making it computationally infeasible to directly search for adversarial patterns across the entire paper. To address this challenge, we first identify the content areas of the paper that attract the most attention from the LLM while generating reviews by feeding the clean paper content into the model and requesting a review.
@@
$=\{m_{1},m_{2},\dots,m_{N}\},$
where $\mathcal{M}$ is the modifiable sequences set for a single paper, $f_{r}(\cdot)$ denotes the review generation by LLM $f$ based on the input content. LCS identifies the longest common sequence matching algorithm.  [^1] LCS can be character-level or word-level depending on the attack method; implementation details could be found in Appendix B., $m$ represents modifiable subsequences, and $N$ is the total number of matched subsequences.
@@
where $\mathcal{P}_{m}$ defines the possible set of perturbations for every modifiable sequence $m$, identified through the transformations specified by the attack method $A$. $\delta$ represents the textual perturbation applied to the original content, and $x^{*}_{adv}$ is the optimal perturbed content found through the search method of the attack. $f_{s}(\cdot)$ represents the sum of the score for all aspects predicted by LLM $f$.
@@
$\text{Total Score Shift}=\sum_{x\in\mathcal{D}}\left(f_{s}(x^{*}_{adv})-f_{s}(x_{clean})\right)$
@@ 3.3 Textual Adversarial Attacks
To examine the robustness of both review generation and score prediction tasks, we investigate a set of representative textual adversarial attack methods. We categorize these methods according to common strategies in textual adversarial attacks. For each strategy, we introduce widely recognized approaches in black-box settings to evaluate the robustness limits of LLMs.
@@
Word-Level Attacks. For word-level adversarial attacks, we implement TextFooler Jin et al. (2020) and BERT-Attack Li et al. (2020) to deceive the LLM. Both attacks focus on replacing words with their synonyms or similar terms in a contextualized embedding space, maintaining semantic similarity while altering the input text.
Sentence-Level Attacks. For the sentence-level attacks, we use StyleAdv Qi et al. (2021a) as representatives for sentence-level attacks. These methods involve transforming sentences in the paper content into specific text styles and modifying stylistic elements without altering the semantic meaning. Such transformations are designed to be imperceptible to human readers but can potentially mislead the LLM.
To conclude, applying these adversarial attack methods enables us to assess the robustness of LLMs in automated peer review. By introducing perturbations at the character, word, and sentence levels, we identify vulnerabilities that compromise the LLM’s reliability. The degree of adversarial changes and examples from different attacks are provided in Appendix C, which offers a visualization of how these perturbations alter the original sentences. Also, the algorithm of textual adversarial attacks can be found in Appendix D.
@@ 4.1 Experiment Settings
LLM Selection and Configuration. We employ both closed- and open-source LLMs for automated peer review. From the closed-source family, we use GPT-4o and its cost-efficient variant GPT-4o-mini, both of which have shown strong performance in text evaluation tasks Robertson (2023); Zhou et al. (2024). For open-source models, we adopt Llama-3.3-70B for its general language capabilities and Mistral-small-3.1 for its competitive performance with GPT-4o-mini, particularly in long-context understanding. Detailed model configurations and prompt templates are provided in Appendix E and F.
Dataset. PeerRead Kang et al. (2018) is one of the earliest publicly released datasets of peer reviews in the NLP community. Following Zhou et al. (2024), we use the ICLR-2017 subset for both review generation and score prediction tasks. This subset contains 427 papers (177 accepted and 255 rejected), with reviews written by experts and aspect-specific scores assigned by annotators. Experimental results on adversarial attack performance and score shift comparison are provided in Appendix G.
AgentReview Jin et al. (2024) presents a large-scale dataset curated through their LLM-based peer review simulation framework. It covers ICLR papers from 2020–2023, sourced from OpenReview, and includes 523 papers (350 rejected, 125 poster, 29 spotlight, and 19 oral).
Since the reviews in both datasets are not manually annotated with aspect tags, we leverage the aspect tagger pre-trained on the ASAP dataset Yuan et al. (2022) to automatically provide aspect tags for human-written reviews.
@@ 4.2 Evaluation Metrics
Review Quality Metrics. To address the research question of how effective LLMs are in generating coherent and insightful reviews compared to human reviewers, we conduct a comparison between LLM-generated reviews and human-annotated reviews, with the latter serving as ground truth. Following the metrics proposed in prior work Yuan et al. (2022); Du et al. (2024); Yu et al. (2024a), we evaluate the quality of automated reviews across multiple dimensions. Our evaluation assesses both the overall quality of LLM-generated reviews and their consistency with human judgments:
(1) Comprehensiveness: For automatically generated reviews, we expect them to cover diverse aspects of the paper. We calculate the Aspect Coverage (ACOV) metric to assess the generated review. In particular, the number of aspects in a predetermined aspect typology that a review has covered can be used to quantify aspect coverage.
[…]
(2) Similarity to human: To evaluate the semantic equivalence between LLM-generated reviews and multiple human-annotated ground truth reviews, we utilize classical statistical methods: ROUGE-1, ROUGE-2, ROUGE-L Lin (2004), and the BertScore metric Zhang et al. (2019). Rather than averaging scores across all human references, we select the maximum score between the LLM-generated review and the human-written ground truths. This approach assumes that closely aligning with every reference is unnecessary; instead, identifying the best match offers a more meaningful assessment of semantic equivalence. Although manually annotated reviews are not free from bias and omissions, human-written evaluations can still be regarded as a practical gold standard.
While these automatic metrics capture surface-level qualities such as fluency and lexical similarity to human-written text, they offer limited insight into deeper dimensions, including coherence, usefulness, and critical insight. To address this gap, we conduct a controlled human evaluation to complement the automatic analysis, as described in Appendix H.1.
@@
In addition to the quantitative results, we conducted a user study to assess the stealthiness of attacks in a realistic peer review setting. Specifically, we evaluated participants’ ability to identify attacked papers and localize the perturbed paragraphs. The detailed settings and full user study results are reported in Appendix H.2.